OpenStudy (anonymous):

What is the product in lowest terms? -8/9 x 5/6= A. –1 B. -13/15 C. -20/27 D. -1/5

OpenStudy (anonymous):

please explain!

OpenStudy (tkhunny):

Well, do it. It might help to shuffle a bit. \(-\dfrac{8}{9}\cdot\dfrac{5}{6} = -\dfrac{8}{6}\cdot\dfrac{5}{9}\). What happened and why might that help?

OpenStudy (anonymous):

I'm not sure @tkhunny

OpenStudy (tkhunny):

8 and 6 have a common factor. Find it and remove it.

OpenStudy (anonymous):

2 is the common factor.

OpenStudy (tkhunny):

That's good. Now reduce the fraction by that common factor.

OpenStudy (anonymous):

4/3 @tkhunny

OpenStudy (tkhunny):

That's the spirit. You are now ready to multiply and you will already have the multiplied form in lowest terms. If you do the simplification FIRST, you won't have to do it later. Go!

OpenStudy (anonymous):

so what do I do next, i'm sorry I am bad at this >.< @tkhunny

OpenStudy (tkhunny):

It is just multiplying fractions. Multiply numerators and multiply denominators. You'll be done in no time. Don't forget the negative sign.